# MeshGraphNet-Transformer: Scalable Mesh-based Learned Simulation for Solid Mechanics

### v1

@@ 2.1 Feature engineering
Each component or part in the physical system is represented as a graph $G(V,E_{M})$, where nodes correspond to mesh vertices $V$ and mesh edges $E_{M}$ encode connectivity. In contact or collision scenarios, the system is modeled as a multi-graph, where initially disjoint component graphs are dynamically connected through contact edges $E_{C}$ when contact occurs. For multi-component bodies, tied constraints are encoded as mesh edges computed using a $k$-nearest neighbor procedure over the assembled components, remaining fixed during simulation.
@@
Positional encoding. We construct a stationary-wave representation over the component’s spatial domain in the undeformed configuration to avoid reliance on absolute positions. Specifically, sinusoidal functions are defined along the length, width, and height of the component [29].
@@ 2.2 Architecture
The Transformer processor performs the global update of node features using two stacked Transformer blocks with physics-attention, see Section 2.3. Before being fed into the Transformer, node attributes are concatenated with positional encodings and projected to match the multi-head attention dimensions. The slicing mechanism of the physics-attention projects the $N$ nodes into $M$ latent physical tokens. By operating on these tokens, the Transformer enables efficient information exchange across the $P$ physical tokens. Then, the de-slicing mechanism projects $P$ back onto $N$ mesh nodes, efficiently completing the global update of the mesh.
@@ 2.3 Physics-attention with eidetic states
We adopt the physics-attention mechanism introduced in Transolver [26] and its extension with eidetic states proposed in Transolver++ [27]. Their proposed physics-attention reduces the quadratic complexity $\mathcal{O}(N^{2})$ of standard self-attention by projecting $N$ node features into a fixed number of latent physical states. This projection not only reduces computational cost but also allows easy handling of varying input sizes and efficient scaling to a large mesh, as the complexity now is $\mathcal{O}(P^{2})$, with $P\ll N$.
To generate the projection, first, a local adaptive temperature mechanism dynamically adjusts the sharpness of the slice weight $\boldsymbol{w}$ between each node and the corresponding physical token,
$\{\boldsymbol{z}_{j}\}_{j=1}^{P}$ $=\left\{\frac{\sum_{i=1}^{N}w_{ij}\boldsymbol{x}_{i}}{\sum_{i=1}^{N}w_{ij}}\right\}_{j=1}^{P},\text{where}\quad\boldsymbol{w}=\mathrm{Softmax}\!\left(\frac{\mathrm{Linear}(\boldsymbol{x})-\log(-\log(\epsilon))}{\tau}\right).$
The weights $\boldsymbol{w}$ are calculated using the slice reparameterization with the Gumbel–Softmax trick [30]. This enables differentiable sampling from the categorical distribution over $P$ physical states $\boldsymbol{z}$, where the Gumble noise is a function of $\epsilon\sim\mathcal{N}(0,1)$. Here, the temperature $\tau$ is learnt as a function of the node features, $\tau=\{\tau_{i}\}_{i=1}^{N}=\{\tau_{0}+\text{Linear}(x_{i})\}_{i=1}^{N}$, to increase expressivity, and $\epsilon$ is random sampled noise. Second, slice reparameterization using the Gumbel–Softmax trick [30], which enables differentiable sampling from the categorical distribution over $M$ physical states $\boldsymbol{z}$. These two techniques, proposed in [27], enable sharper weighting and prevent tokens from collapsing, providing a better latent representation of the physical token states.
Here, physics-attention is defined as a multi-head attention mechanism [29] over the $P$ physical tokens. Each token is linearly projected into Query ($Q$), Key ($K$), and Value ($V$), with dimensionality $c$. The scaled dot product between the Q{Q} and the $K$ produces the attention weights, which are normalised with a $\mathrm{Softmax}$ function. Afterwards, the attention weights are projected onto the $V$ to generate the updated $\boldsymbol{z^{\prime}}$ physical $P$ tokens,
@@ 3 Numerical Examples
We consider Lagrangian mesh-based models undergoing large deformations in two distinct regimes: a dynamic impact scenario (pi-beam benchmark) and a quasi-static benchmark problem coined as deforming plate. An overview of both datasets is provided in Table 1. The pi-beam dataset is described in detail in Section 3.1. The deforming plate dataset is a widely used public benchmark introduced by Pfaff et al. [5] for evaluating surrogate models in solid mechanics.
@@ 3.1 Pi-beam benchmark
The geometry is discretized using shell elements and a total 16,193 nodes per simulation, where 15,641 belong to the deformable structure and 552 to the rigid obstacle, see Figure 2(b). The learning task consists of predicting the full nonlinear dynamic response of the system, including contact, collisions, and subsequent spring-back. Node state variables $\boldsymbol{s}_{i}$ include thickness $k_{i}$, positions $\boldsymbol{x}_{i}$, velocity $\boldsymbol{v}_{i}$, and the plastic hardening variable $\alpha_{i}$, which captures the evolving material properties during loading cases (see Table 1). The type of node (rigid, deformable, clamped, …) is also stored as a one-hot encoder vector $\boldsymbol{n}_{i}$. From each simulation we pruned 100 snapshots, corresponding to a total physical time of 0.15 s, or 0.0015 s per considered step.
The dataset is generated using the Pam Crash software (ESI Group, Keysight Technologies), which employs an explicit finite element method (FEM) solver with a temporal resolution of $6.3\times 10^{-4}\mathrm{s}$. The material is assumed to be steel, elastoplastic, with power-law hardening and Young’s modulus $E=210000$ MPa. One single FE simulation took roughly ten minutes employing four processors. The resulting dataset consists of 28 trajectories in total: 18 for training and validation, and 10 for testing.
@@
Deforming Plate 400 1,250 Teth. $\boldsymbol{u}_{i}^{\text{imp}},\,\boldsymbol{n}_{i}$ $\boldsymbol{u}_{i}^{k+1},\,\sigma_{\text{vM},i}^{k+1}$
@@
3.2 Deforming plate
While the pi-beam example is intended to check the robustness of the proposed method in the face of changes in the simulation parameter values (in this case, the thicknesses of the different components of the structure), the deforming plate example aims to investigate robustness in the face of changes in the geometry, boundary conditions, and mesh of the , see [4].
@@ 4 Experiments
We evaluate the model’s performance by evaluating the root mean squared error on 1-step predictions (RMSE-1) and the full-trajectory (RMSE-all). In addition, we report physical consistency metrics and qualitative visualizations to further analyze the simulated dynamics.
RMSE-all is computed as the root-mean-square error between predictions and FEM solutions, taken as ground truth, per variable. For each error, we calculate the mean squared error across all feature variables, nodes, time steps, and trajectories, then take the square root (Described in [5], Appendix A.5.2 ). RMSE-1 is computed without autoregressive rollout: predictions always start from the ground-truth state to prevent error accumulation. Relative RMSE (R-RMSE) follows the same computation, with each error normalized by the infinity norm of the corresponding trajectory (i.e., its maximum magnitude).
4.1 Pi-Beam dataset
Model MPNN Heads Tokens Hidden #Params
@@ 4.1 Pi-Beam dataset
On the pi-beam dataset, a standard MGN would require more than 175 message-passing iterations to cover the entire component. This lower bound is dictated by the criterion established in [18]. Using fewer iterations leads to under-reaching, preventing the network from capturing the relevant physics and resulting in inaccurate predictions. Consequently, this architecture becomes intractable for this example, and even if made tractable, it would be slow due to the large number of required iterations. This scenario is common in industrial-scale simulations like this one. Nevertheless, we include MGN as a baseline using its original configuration. Details about this architecture can be found at Table 2.
Table 3 reports a comparison of absolute and relative errors. As expected, MGN-T outperforms MGN across all variables and metrics, as MGN suffers from under-reach due to insufficient message-passing iterations. In contrast, MGN-T successfully overcomes this limitation, maintaining the relative errors below 5%, as shown in Figure 3.
Figure 4(a) illustrates the prediction of a full trajectory and how the nodes that undergo plasticity evolve, highlighting that these are highly localized. Figure 4(b) shows the velocity magnitude at each time step for consecutive predictions, starting from the beginning of the trajectory. Two key observations: MGN-T demonstrates outstanding time resolution capabilities compared to the FEM method, which cannot operate at this temporal resolution. Additionally, MGN-T accurately predicts buckling, particularly high on the top left side of the component.
@@
$\boldsymbol{q}$(m) $\boldsymbol{v}$(m/s) $\alpha\times 10^{-2}$(m/m) $\boldsymbol{q}$(%) $\boldsymbol{v}$(%) $\alpha$(%)
@@
Figure 3: Error evolution rollout comparison between FEM ground truth and MGN-T predictions for the pi-beam benchmark. Line represents the mean and shade the standard error accross trajectories. Top: RMSE. Bottom: Relative RMSE.
@@
Figure 4: Prediction MGN-T Pi-Beam dataset: MGN-T prediction on the for a dynamic impact against an obstacle. a) Temporal evolution of hardening variable $\alpha$. b) Temporal evolution of velocity magnitude $\|\boldsymbol{v}\|$ for consecutive temporal predictions, from beginning of trajectory.
@@
Plasticity evolution and physical consistency: We evaluate physical consistency by analyzing the evolution of the plastic hardening variable $\alpha$, which, according to yield criteria, should increase monotonically. As a proxy for measuring the level of plastification, we visualise the sum of hardening across nodes at each temporal increment. Similarly, we perform the same proxy for the velocity magnitude. Figure 5 compares the predicted and ground-truth plastic hardening and system kinetic energy over time. MGN-T closely follows the FEM solution, preserving monotonic hardening throughout the simulation, in agreement with plasticity theory. The same can be observed for the velocity, which decreases as the system dissipates energy during plastification. These results indicate that MGN-T not only achieves low numerical error but also respects key physical constraints of plasticity and the corresponding relationships to other state variables.
@@
Figure 5: Physical consistency Pi-Beam predictions. Top: Accumulated plastic hardening. Bottom: Kinetic Energy magnitude. Each column is one different test trajectory, for ground truth, MGN-T and MGN.
Physics-attention visualization and interpretability: To analyze how MGN-T captures long-range physical interactions, we visualize the learned physics-attention mechanisms in the Transformer processor. The attention maps illustrate how information from the full mesh is aggregated into a small set of $P$ physics tokens and subsequently redistributed back to individual nodes.
@@
From our perspective, these physics tokens can be interpreted as a dynamic hierarchical representation of the system. Unlike classical hierarchical GNNs, where the coarse-to-fine mapping is fixed by a predefined mesh coarsening strategy and determined solely by the spatial domain, the tokenization in MGN-T is learned and continuously adapts based on the current latent nodal states and their encoded spatial context. This allows the model to construct an implicit hierarchy that evolves throughout the simulation, responding jointly to changes in state variables and geometry.
While directly interpreting such learned representations remains challenging, visualizing the attention weights provides valuable qualitative insight into what the model has learnt under the hood. We observe that attention patterns vary across time. They concentrate on physically relevant regions, such as highly deformed and plastified regions (see Figure 6(a)), impact surfaces (see Figure 6(d)), or broader regions, such as the lower part of the component (see Figure 6(c)). These patterns shift as the simulation evolves. This indicates that the model dynamically reallocates global attention in response to changing physical state variables (see Figure 6).
@@ 4.2 Deforming plate
MGN [5] 0.26 15.1 2.0M $\checkmark$ $\times$ $\checkmark$ $\checkmark$
@@
MGN-T (ours) 0.11 3.21 0.5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 5: Performance MGN-T model on Deforming Plate dataset. RMSE and relative RMSE for $\boldsymbol{q}$ and $\sigma_{\rm VM}$.
@@
For completeness, we additionally report a comparison table of state-of-the-art baselines that benchmarked their models, using the metrics reported in the original publications. All the compared baselines are mesh-based methods. In this table, we present the single-step and rollout errors (RMSE-1 and RMSE-All, respectively), together with the total number of trainable parameters for each model (Table 4). This allows for a direct overview of the reported performance–complexity trade-offs across methods. For qualitative visualization of the results, see Figure 7.
Our proposed method achieves competitive performance with state-of-the-art approaches, ranking second in predictive accuracy. Importantly, this result is obtained using only a fraction of the parameters required by other methods. Unlike hierarchical or segment-based models, MGN-T operates directly at the node level and does not require any preprocessing steps such as mesh–graph segmentation, mesh coarsening, or prior domain knowledge. As a result, it offers a more straightforward, end-to-end training pipeline while demonstrating strong off-the-shelf performance, as shown on the right side of Table 4.
@@
MGN-T achieves competitive performance with state-of-the-art approaches, ranking second in accuracy (but note that it provides a relative error value that is judged sufficient for the vast majority of applications, on the order of $0.3\%$). Importantly, this result is obtained using only a fraction of the parameters required by other methods. The proposed MGN-T is the only non-hierarchical and coarsening-free method that achieves this level of performance. The Transformer provides an efficient way to perform a global update to prevent under-reaching, which is the main reason for failure of MGN. HCMT is the only method from the baseline that does not rely on MPNNs but instead uses mesh transformer blocks. After the Contact Transformer (CT) process, the collision between the actuator and the plate, stacked blocks of Hierarchical Mesh Transformers (HMT), followed by pooling layers, coarsens the graph. For upsampling, a similar pipeline is followed, using unpooling layers to update the nodes at the finest level. The level of coarsening is controlled by the number of stacked blocks and it is end-to-end trainable.
@@
Hierarchical approaches, such as BSMS-GNN construct a hierarchical graph by bi-stride pooling, which groups nodes every other topological depth. The coarse levels are computed prior to training and are non-adaptive. This depth is estimated from the geodetic distance of the mesh using Breadth-First-Search (BFS) from a selected node. This technique is an extension of bipartitioning, which guarantees 2-CC for Directed Acyclic Graphs (DAGs), but, while well-suited to DAGs, might fail to provide the appropriate hierarchy for sophisticated meshes. Here, downsampling and upsampling layers, which connect different coarsening levels, are non-parametric weighted aggregations. On the other hand, EvoMesh relies on differentiable multi-scale graph construction, which has a learnable probability distribution over whether to retain or discard a node at the next coarser graph level. When constructing the multi-level connectivity, the edges of the original mesh are used. Additionally, it incorporates $k$-hop edges to ensure no disconnected graph partitions, which would limit the message-passing information flow. For communication across coarser levels, reduced and expand layers are learned using the weights estimated by their proposed anisotropic message passing layers. This method provides a learnable hierarchy construction and coarsening that adapt to latent nodal states, at the cost of increasing the model’s trainable parameters.
@@
As mentioned above, MGN-T is the lightest and best-performing method that does not require hierarchy or coarsening and can be trained end-to-end, without under-reaching despite topology changes and sophisticated meshes.

### v3

@@ 2.1 Feature engineering
Each component or part in the physical system is represented as a graph $G(V,E_{M})$, where nodes correspond to mesh vertices $V$ and mesh edges $E_{M}$ encode connectivity. In contact or collision scenarios, the system is modeled as a multi-graph, where initially disjoint component graphs are dynamically connected through contact edges $E_{C}$ when contact occurs. For multi-component bodies, tied constraints are encoded as mesh edges computed using a $k$-nearest neighbour procedure over the assembled components, remaining fixed during simulation.
@@
Positional encoding: We construct a stationary-wave representation over the component’s spatial domain in the undeformed configuration to avoid reliance on absolute positions. Specifically, sinusoidal functions are defined along the length, width, and height of the component [16].
@@ 2.2 Architecture
The Transformer processor performs the global update of node features using two stacked Transformer blocks with physics-attention, see Section 2.3. Before being fed into the Transformer, node attributes are concatenated with positional encodings and projected to match the multi-head attention dimensions. The slicing mechanism of the physics-attention projects the $N$ nodes into $P$ latent physical tokens. By operating on these tokens, the Transformer enables efficient information exchange across these. Then, the de-slicing mechanism projects $P$ back onto $N$ mesh nodes, efficiently completing the global update of the mesh.
@@ 2.3 Physics-attention with eidetic states
We adopt the physics-attention mechanism introduced in Transolver [29] and its extension with eidetic states proposed in Transolver++ [28]. Their proposed physics-attention reduces the quadratic complexity $\mathcal{O}(N^{2})$ of standard self-attention by projecting $N$ node features into $P$ physical tokens, as $\boldsymbol{z}$ eidetic states. This projection not only reduces computational cost but also allows easy handling of varying input sizes and efficient scaling to a large mesh, as the complexity now is $\mathcal{O}(P^{2})$, with $P\ll N$.
To generate the projection, first, a local adaptive temperature mechanism dynamically adjusts the sharpness of the slice weight $\boldsymbol{w}$ between each node and the corresponding physical token. Here, the temperature $\tau$ is learnt from the node features, $\tau=\{\tau_{i}\}_{i=1}^{N}=\{\tau_{0}+\text{Linear}(x_{i})\}_{i=1}^{N}$, to increase expressivity to learn relationships between nodes and physical tokens,
$\{\boldsymbol{z}_{j}\}_{j=1}^{P}=\left\{\frac{\sum_{i=1}^{N}w_{ij}\boldsymbol{x}_{i}}{\sum_{i=1}^{N}w_{ij}}\right\}_{j=1}^{P},$
where
$\boldsymbol{w}=\mathrm{Softmax}\!\Bigg(\frac{\mathrm{Linear}(\boldsymbol{x})-\log(-\log(\epsilon))}{\tau}\Bigg).$
Second, the slice reparameterization mechanism with the Gumbel–Softmax trick [31] used to calcualte the weights $\boldsymbol{w}$. This enables differentiable sampling from the categorical distribution over physical states $\boldsymbol{z}$, where the Gumble noise is a function of $\epsilon\sim\mathcal{N}(0,1)$. These two techniques, proposed in [28], enable sharper weighting and prevent tokens from collapsing, providing a better latent representation of the physical token states.
Physics-attention is defined as a multi-head attention mechanism [16] over the $P$ physical tokens, where $P$ is a hyperparameter. Each token is linearly projected into Query ($Q$), Key ($K$), and Value ($V$), with dimensionality $c$. The scaled dot product between the Q{Q} and the $K$ produces the attention weights, which are normalised with a $\mathrm{Softmax}$ function. Afterwards, the attention weights are projected onto the $V$ to generate the updated $\boldsymbol{z^{\prime}}$ physical $P$ tokens,
@@ 3 Numerical Examples
We consider Lagrangian mesh-based models undergoing large deformations in two distinct regimes: a dynamic impact scenario (Pi-beam benchmark) and a quasi-static scenario problem coined as Deforming plate, a widely used public benchmark introduced by Pfaff et al. [18] for evaluating surrogate models in solid mechanics. An overview of both datasets is provided in Table 1. The Pi-beam dataset is described in detail in Section 3.1 and Deforming plate dataset in Section 3.2.
@@ 3.1 Pi-beam benchmark
The geometry is discretized using shell elements and a total 16,193 nodes per simulation, where 15,641 belong to the deformable structure and 552 to the rigid obstacle, see Figure 2(b). The learning task consists of predicting the full nonlinear dynamic response of the system, including contact, collisions, and subsequent spring-back. Node state variables $\boldsymbol{s}_{i}$ include thickness $k_{i}$, positions $\boldsymbol{x}_{i}$, velocity $\boldsymbol{v}_{i}$, and the plastic hardening variable $\alpha_{i}$, which captures the evolving material properties during loading cases (see Table 1). The type of node (rigid, deformable, clamped, …) is also stored as a one-hot encoder vector $\boldsymbol{n}_{i}$. From each simulation we pruned 100 snapshots, corresponding to a total physical time of 15 ms, or 1.5 ms per considered step.
The dataset is generated using the Pam Crash software (ESI Group, Keysight Technologies), which employs an explicit finite element method (FEM) solver with a temporal resolution of $6.3\times 10^{-4}\mathrm{ms}$. The material is assumed to be steel, elastoplastic, with power-law hardening and Young’s modulus $E=210000$ MPa. One single FE simulation took roughly ten minutes employing four processors and mass scaling. The resulting dataset consists of 28 trajectories in total: 18 for training and validation, and 10 for testing.
@@
Def. Plate 400 1,250 Teth. $\boldsymbol{u}_{i}^{\text{imp}},\,\boldsymbol{n}_{i}$ $\boldsymbol{u}_{i}^{k+1},\,\sigma_{\text{vM},i}^{k+1}$
@@
3.2 Deforming plate benchmark
While the pi-beam example is intended to check the robustness of the proposed method in the face of changes in the simulation parameter values (in this case, the thicknesses of the different components of the structure), the deforming plate example aims to investigate robustness in the face of changes in the geometry, boundary conditions, and mesh, see [17].
@@ 4 Experiments
We evaluate the model’s performance by means of the root mean squared error on 1-step predictions (RMSE-1) and the full-trajectory (RMSE-all). In addition, we report physical consistency metrics and qualitative visualizations to further analyze the simulated dynamics.
RMSE is computed as the root-mean-square error between predictions and FEM solutions, taken as ground truth, per variable. For each error, we calculate the mean squared error across all feature variables, nodes, time steps, and trajectories, then take the square root (Described in [18], Appendix A.5.2 ). RMSE-1 is computed without autoregressive rollout: predictions always start from the ground-truth state to prevent error accumulation. Relative RMSE (R-RMSE) follows the same computation, with each error normalized by the infinity norm of the corresponding trajectory (i.e., its maximum magnitude). From now on RMSE and RMSE-all are equivalente.
4.1 Pi-beam
Model MPNN Heads Tokens ($P$) Hidden #Params
@@ 4.1 Pi-beam
On the pi-beam dataset, a standard MGN would require more than 175 message-passing iterations to cover the entire component. This lower bound is dictated by the criterion established in [22]. Using fewer iterations leads to under-reaching, preventing the network from capturing the relevant physics and resulting in inaccurate predictions. Consequently, this architecture becomes intractable for this example, and even if made tractable, it would be slow due to the large number of required iterations. This scenario is common in industrial-scale simulations like this one. Nevertheless, we include MGN as a baseline using its original configuration of 15-layer MPNNs proposed in [18]. Details about this architecture can be found at Table 2.
Table 3 reports a comparison of absolute and relative errors. As expected, MGN-T outperforms MGN across all variables and metrics, as MGN suffers from under-reach due to insufficient message-passing iterations. In contrast, MGN-T successfully overcomes this limitation, maintaining the relative errors way below 5%, as shown in Figure 3.
Figure 4(a) illustrates the prediction of a full trajectory and how the nodes that undergo plasticity evolve, highlighting that these are highly localized. Figure 4(b) shows the velocity magnitude at each time step for consecutive predictions, starting from the beginning of the trajectory. MGN-T accurately predicts buckling, particularly high on the top left side of the component, at a temporal resolution where FEM cannot operate, demonstrating its outstanding capability to handle large temporal steps.
@@
$\boldsymbol{q}$($m$) $\boldsymbol{v}$(mm/ms) $\alpha\times 10^{-2}$(mm/$m$) $\boldsymbol{q}$(%) $\boldsymbol{v}$(%) $\alpha$(%)
@@
Figure 3: Error evolution rollout comparison between FEM ground truth and MGN-T predictions for the pi-beam benchmark. Line represents the mean and shade the standard error accross trajectories. Top: RMSE. Bottom: R-RMSE.
@@
Figure 4: Prediction MGN-T Pi-Beam dataset: MGN-T prediction for a dynamic impact against an obstacle. a) Temporal evolution of hardening variable $\alpha$. b) Temporal evolution of velocity magnitude $\|\boldsymbol{v}\|$ for consecutive temporal predictions.
@@
Plasticity evolution and physical consistency: We evaluate physical consistency by analyzing the evolution of the plastic hardening variable $\alpha$, which, according to yield criteria, should increase monotonically. As a proxy for measuring the level of plastification, we visualise the sum of hardening across nodes at each temporal increment. Similarly, for the velocity we calculate the kinetic energy. Figure 5 compares the predicted and ground-truth plastic hardening and system kinetic energy over time. MGN-T closely follows the FEM solution, preserving monotonic hardening throughout the simulation, in agreement with plasticity theory. The same consistency is observed for the kinetic energy, which decreases as the system dissipates energy during plastification and gathers potential elastic energy, then recovered as kinetic energy during springback. These results indicate that MGN-T respects key physical constraints of plasticity and the corresponding relationships with other state variables.
@@
Figure 5: Physical consistency of the pi-beam predictions. Top: Accumulated plastic hardening. Bottom: Kinetic Energy magnitude. Each column represents a different test trajectory, for ground truth, MGN-T and MGN.
@@
Physics-attention visualization and interpretability: While directly interpreting the learned token representations remains challenging, visualizing the slice weights $\boldsymbol{w}$ provides valuable qualitative insight into what the model has learnt under the hood (see Figure 6). The weights $\boldsymbol{w}$ illustrate how information from the full mesh is aggregated into each $P$ token. From our perspective, these physics tokens can be interpreted as a dynamic hierarchical representation of the system. Unlike classical hierarchical GNNs, where the coarse-to-fine mapping is fixed by a predefined mesh coarsening strategy and determined solely by the spatial domain, the tokenization in MGN-T is learned and adapts based on the current latent nodal states and their encoded spatial context. This allows the model to construct an implicit hierarchy that evolves throughout the simulation, responding to changes in state variables and geometry. The weights concentrate on physically relevant regions, such as highly deformed and plastified regions (see Figure 6(a)), impact surfaces (see Figure 6(d)), or broader regions, such as the lower part of the component (see Figure 6(c)).
@@ 4.2 Deforming plate
MGN [18] 0.25 15.1 2.0M $\checkmark$ $\times$ $\checkmark$ $\checkmark$
@@
MGN-T (ours) 0.10 3.2 0.5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
[…]
Table 5: Performance of the MGN-T model on the Deforming Plate dataset. RMSE and relative RMSE for $\boldsymbol{q}$ and $\sigma_{\rm VM}$, reported for multi-step and 1-step predictions.
@@
For completeness, we additionally report a comparison table of state-of-the-art baselines benchmarked on the dataset. All the compared baselines are mesh-based methods. In this table, we present the single-step and rollout errors (RMSE-1 and RMSE-All, respectively), together with the total number of trainable parameters for each model (Table 4). This allows for a direct overview of the reported performance–complexity trade-offs across methods. For qualitative visualization of the results, see Figure 7.
@@
Our proposed method achieves competitive performance with state-of-the-art approaches, ranking second in accuracy (but note that it provides a relative error value that is judged sufficient for the vast majority of applications, on the order of $0.6\%$; see Table 5). Importantly, this result is obtained using only a fraction of the parameters required by other methods. The proposed MGN-T is the only non-hierarchical and coarsening-free method that achieves this level of performance. The Transformer provides an efficient way to perform a global update to prevent under-reaching, which is the main reason for failure of MGN.
@@
HCMT is the only method from the baseline that does not rely on MPNNs but instead uses mesh transformer blocks. Here, stacked blocks of Hierarchical Mesh Transformers (HMT), followed by pooling layers, coarsens the graph. Hierarchical approaches, such as BSMS-GNN construct a hierarchical graph by bi-stride pooling, which groups nodes every other topological depth. The coarse levels are computed prior to training and are non-adaptive. This depth is estimated from the geodetic distance of the mesh using Breadth-First-Search (BFS) from a selected node. This technique is an extension of bipartitioning, which guarantees 2-CC for Directed Acyclic Graphs (DAGs), but, while well-suited to DAGs, might fail to provide the appropriate hierarchy for sophisticated meshes. Here, downsampling and upsampling layers, which connect different coarsening levels, are non-parametric weighted aggregations. On the other hand, EvoMesh relies on differentiable multi-scale graph construction, which has a learnable probability distribution over whether to retain or discard a node at the next coarser graph level. When constructing the multi-level connectivity, the edges of the original mesh are used. Additionally, it incorporates $k$-hop edges to ensure no disconnected graph partitions, which would limit the message-passing information flow. For communication across coarser levels, reduced and expand layers are learned using the weights estimated by their proposed anisotropic message passing layers. This method provides a learnable hierarchy construction and coarsening that adapt to latent nodal states, at the cost of increasing the model’s trainable parameters. Nevertheless, despite learnable coarsening being an advantage, it seems like the heuristics employed might have limited the performance of these methods.
@@
As mentioned above, MGN-T is the lightest and best-performing method that does not require hierarchy or coarsening. It can be trained end-to-end, offering a more straightforward training pipeline while demonstrating strong off-the-shelf performance.